Determine the interest payment for the following three bonds. (Assume a $1,000 par value.) (Round your answers to 2 decimal places.)

4.75 percent coupon corporate bond (paid semiannally) __________

5.40 percent coupon treasury note _______

Corporate zero-coupon bond maturing in ten years ______
Answer #1

Computation of Interest payment for the bonds

Par value * coupon rate * N / 12

1) 4.75% (paid semi-annually)

= $1,000 * 4.75% * 6/12 = $23.75

2) 5.40%

$1,000*5.40% = $54.

for 6 months it is $27

3) Zero-coupon bond has no interest

$1,000 * 0% = $0